### v1

@@ Acknowledgements.
We acknowledge G. Mattoni, H. Matsuki, and C. Sow for their technical supports. We also acknowledge L. Balents and J. N. Hausman for useful comments and discussions. This work was supported by JST-CREST (No. JPMJCR18T1), Grant-in-Aid for Scientific Research (JSPS KAKENHI Grant Numbers JP22H01168, JP21H04435 and JP19H01845), Grant-in-Aid for JSPS Research Fellow (No: JP18J20058), and KEK-PF (Proposal number: 2024S2-001). T. Osumi thanks GP-Spin and A.H. thanks GP-Spin and JSPS for financial support.References
- Bychkov and Rashba [1984] Y. A. Bychkov and E. I. Rashba, Properties of a 2D electron gas with lifted spectral degeneracy, P. Zh. Eksp. Teor. Fiz. 39, 66 (1984).
- Dresselhaus [1955] G. Dresselhaus, Spin-Orbit Coupling Effects in Zinc Blende Structures, Phys. Rev. 100, 580 (1955).
- Murakami et al. [2003] S. Murakami, N. Nagaosa, and S.-C. Zhang, Dissipationless Quantum Spin Current at Room Temperature, Science 301, 1348 (2003).
- Sinova et al. [2004] J. Sinova, D. Culcer, Q. Niu, N. A. Sinitsyn, T. Jungwirth, and A. H. MacDonald, Universal Intrinsic Spin Hall Effect, Physical Review Letters 92, 126603 (2004).
- Kato et al. [2004] Y. K. Kato, R. C. Myers, A. C. Gossard, and D. D. Awschalom, Observation of the spin hall effect in semiconductors, Science 306, 1910 (2004).
- Sinova et al. [2015] J. Sinova, S. O. Valenzuela, J. Wunderlich, C. H. Back, and T. Jungwirth, Spin Hall effects, Reviews of Modern Physics 87, 1213 (2015).
- Sigrist [2009] M. Sigrist, Introduction to unconventional superconductivity in non‐centrosymmetric metals, AIP Conference Proceedings 1162, 55 (2009).
- Yip [2014] S. Yip, Noncentrosymmetric Superconductors, Annual Review of Condensed Matter Physics 5, 15 (2014).
- Daido et al. [2022] A. Daido, Y. Ikeda, and Y. Yanase, Intrinsic Superconducting Diode Effect, Physical Review Letters 128, 037001 (2022).
- Yuan and Fu [2022] N. F. Q. Yuan and L. Fu, Supercurrent diode effect and finite-momentum superconductors, Proceedings of the National Academy of Sciences 119, e 2119548119 (2022).
- Ando et al. [2020] F. Ando, Y. Miyasaka, T. Li, J. Ishizuka, T. Arakawa, Y. Shiota, T. Moriyama, Y. Yanase, and T. Ono, Observation of superconducting diode effect, Nature 584, 373 (2020).
- Ahn et al. [2019] K. H. Ahn, A. Hariki, K. W. Lee, and J. Kuneš, Antiferromagnetism in RuO2 as -wave Pomeranchuk instability, Physical Review B 99, 184432 (2019).
- Naka et al. [2019] M. Naka, S. Hayami, H. Kusunose, Y. Yanagi, Y. Motome, and H. Seo, Spin current generation in organic antiferromagnets, Nature Communications 10, 4305 (2019).
- Hayami et al. [2019] S. Hayami, Y. Yanagi, and H. Kusunose, Momentum-Dependent Spin Splitting by Collinear Antiferromagnetic Ordering, Journal of the Physical Society of Japan 88, 123702 (2019).
- Šmejkal et al. [2020] L. Šmejkal, R. González-Hernández, T. Jungwirth, and J. Sinova, Crystal time-reversal symmetry breaking and spontaneous Hall effect in collinear antiferromagnets, Science Advances 6, eaaz8809 (2020).
- Yuan et al. [2020] L.-D. Yuan, Z. Wang, J.-W. Luo, E. I. Rashba, and A. Zunger, Giant momentum-dependent spin splitting in centrosymmetric low-Z antiferromagnets, Physical Review B 102, 014422 (2020).
- Yuan et al. [2021] L.-D. Yuan, Z. Wang, J.-W. Luo, and A. Zunger, Prediction of low-Z collinear and noncollinear antiferromagnetic compounds having momentum-dependent spin splitting even without spin-orbit coupling, Physical Review Materials 5, 014409 (2021).
- Ma et al. [2021] H.-Y. Ma, M. Hu, N. Li, J. Liu, W. Yao, J.-F. Jia, and J. Liu, Multifunctional antiferromagnetic materials with giant piezomagnetism and noncollinear spin current, Nature Communications 12, 2846 (2021).
- Šmejkal et al. [2022a] L. Šmejkal, J. Sinova, and T. Jungwirth, Beyond Conventional Ferromagnetism and Antiferromagnetism: A Phase with Nonrelativistic Spin and Crystal Rotation Symmetry, Physical Review X 12, 031042 (2022a).
- Šmejkal et al. [2022b] L. Šmejkal, J. Sinova, and T. Jungwirth, Emerging Research Landscape of Altermagnetism, Physical Review X 12, 040501 (2022b).
- González-Hernández et al. [2021] R. González-Hernández, L. Šmejkal, K. Výborný, Y. Yahagi, J. Sinova, T. Jungwirth, and J. Železný, Efficient Electrical Spin Splitter Based on Nonrelativistic Collinear Antiferromagnetism, Physical Review Letters 126, 127701 (2021).
- Feng et al. [2022] Z. Feng, X. Zhou, L. Šmejkal, L. Wu, Z. Zhu, H. Guo, R. González-Hernández, X. Wang, H. Yan, P. Qin, X. Zhang, H. Wu, H. Chen, Z. Meng, L. Liu, Z. Xia, J. Sinova, T. Jungwirth, and Z. Liu, An anomalous Hall effect in altermagnetic ruthenium dioxide, Nature Electronics 5, 735 (2022).
- Bai et al. [2022] H. Bai, L. Han, X. Y. Feng, Y. J. Zhou, R. X. Su, Q. Wang, L. Y. Liao, W. X. Zhu, X. Z. Chen, F. Pan, X. L. Fan, and C. Song, Observation of Spin Splitting Torque in a Collinear Antiferromagnet RuO2, Physical Review Letters 128, 197202 (2022).
- Karube et al. [2022] S. Karube, T. Tanaka, D. Sugawara, N. Kadoguchi, M. Kohda, and J. Nitta, Observation of Spin-Splitter Torque in Collinear Antiferromagnetic RuO2, Physical Review Letters 129, 137201 (2022).
- Bose et al. [2022] A. Bose, N. J. Schreiber, R. Jain, D.-F. Shao, H. P. Nair, J. Sun, X. S. Zhang, D. A. Muller, E. Y. Tsymbal, D. G. Schlom, and D. C. Ralph, Tilted spin current generated by the collinear antiferromagnet ruthenium dioxide, Nature Electronics 5, 267 (2022).
- Bai et al. [2023] H. Bai, Y. C. Zhang, Y. J. Zhou, P. Chen, C. H. Wan, L. Han, W. X. Zhu, S. X. Liang, Y. C. Su, X. F. Han, F. Pan, and C. Song, Efficient Spin-to-Charge Conversion via Altermagnetic Spin Splitting Effect in Antiferromagnet RuO2, Physical Review Letters 130, 216701 (2023).
- Tschirner et al. [2023] T. Tschirner, P. Keßler, R. D. G. Betancourt, T. Kotte, D. Kriegner, B. Büchner, J. Dufouleur, M. Kamp, V. Jovic, L. Smejkal, J. Sinova, R. Claessen, T. Jungwirth, S. Moser, H. Reichlova, and L. Veyrat, Saturation of the anomalous Hall effect at high magnetic fields in altermagnetic RuO2, APL Materials 11, 101103 (2023).
- Wang et al. [2023] M. Wang, K. Tanaka, S. Sakai, Z. Wang, K. Deng, Y. Lyu, C. Li, D. Tian, S. Shen, N. Ogawa, N. Kanazawa, P. Yu, R. Arita, and F. Kagawa, Emergent zero-field anomalous Hall effect in a reconstructed rutile antiferromagnetic metal, Nature Communications 14, 8240 (2023).
- Guo et al. [2024] Y. Guo, J. Zhang, Z. Zhu, Y. yuan Jiang, L. Jiang, C. Wu, J. Dong, X. Xu, W. He, B. He, Z. Huang, L. Du, G. Zhang, K. Wu, X. Han, D. fu Shao, G. Yu, and H. Wu, Direct and Inverse Spin Splitting Effects in Altermagnetic RuO2, Advanced Science 11, 2400967 (2024).
- Wang et al. [2024] Z. Wang, Z. Li, L. Sun, Z. Zhang, K. He, H. Niu, J. Cheng, M. Yang, X. Yang, G. Chen, Z. Yuan, H. Ding, and B. Miao, Inverse Spin Hall Effect Dominated Spin-Charge Conversion in (101) and (110)-Oriented RuO2 Films, Physical Review Letters 133, 046701 (2024).
- Jeong et al. [2024] S. G. Jeong, I. H. Choi, S. Nair, L. Buiarelli, B. Pourbahari, J. Y. Oh, N. Bassim, A. Seo, W. S. Choi, R. M. Fernandes, T. Birol, L. Zhao, J. S. Lee, and B. Jalan, Altermagnetic Polar Metallic phase in Ultra-Thin Epitaxially-Strained RuO2 Films, arXiv:2405.05838 (2024).
- Berlijn et al. [2017] T. Berlijn, P. Snijders, O. Delaire, H.-D. Zhou, T. Maier, H.-B. Cao, S.-X. Chi, M. Matsuda, Y. Wang, M. Koehler, P. Kent, and H. Weitering, Itinerant Antiferromagnetism in RuO2, Physical Review Letters 118, 077201 (2017).
- Zhu et al. [2019] Z. H. Zhu, J. Strempfer, R. R. Rao, C. A. Occhialini, J. Pelliciari, Y. Choi, T. Kawaguchi, H. You, J. F. Mitchell, Y. Shao-Horn, and R. Comin, Anomalous Antiferromagnetism in Metallic RuO2 Determined by Resonant X-ray Scattering, Physical Review Letters 122, 17202 (2019).
- Gregory et al. [2022] B. Z. Gregory, J. Strempfer, D. Weinstock, J. P. Ruf, Y. Sun, H. Nair, N. J. Schreiber, D. G. Schlom, K. M. Shen, and A. Singer, Strain-induced orbital-energy shift in antiferromagnetic RuO2 revealed by resonant elastic x-ray scattering, Physical Review B 106, 195135 (2022).
- Zhou et al. [2024] X. Zhou, W. Feng, R.-W. Zhang, L. Šmejkal, J. Sinova, Y. Mokrousov, and Y. Yao, Crystal Thermal Transport in Altermagnetic RuO2, Physical Review Letters 132, 056701 (2024).
- Shao et al. [2021] D. F. Shao, S. H. Zhang, M. Li, C. B. Eom, and E. Y. Tsymbal, Spin-neutral currents for spintronics, Nature Communications 12, 7061 (2021).
- Shao et al. [2023] D. F. Shao, Y. Y. Jiang, J. Ding, S. H. Zhang, Z. A. Wang, R. C. Xiao, G. Gurung, W. J. Lu, Y. P. Sun, and E. Y. Tsymbal, Néel Spin Currents in Antiferromagnets, Physical Review Letters 130, 1 (2023).
- Liu et al. [2023] Y. Liu, H. Bai, Y. Song, Z. Ji, S. Lou, Z. Zhang, C. Song, and Q. Jin, Inverse Altermagnetic Spin Splitting Effect-Induced Terahertz Emission in RuO2, Advanced Optical Materials 11, 2300177 (2023).
- Zhang et al. [2023] S. Zhang, Y. Cui, S. Wang, H. Chen, Y. Liu, W. Qin, T. Guan, C. Tian, Z. Yuan, L. Zhou, Y. Wu, and Z. Tao, Nonrelativistic and nonmagnetic terahertz-wave generation via ultrafast current control in anisotropic conductive heterostructures, Advanced Photonics 5, 056006 (2023).
- Over et al. [2000] H. Over, Y. D. Kim, A. P. Seitsonen, S. Wendt, E. Lundgren, M. Schmid, P. Varga, A. Morgante, and G. Ertl, Atomic-Scale Structure and Catalytic Reactivity of the RuO2 (110) Surface, Science 287, 1474 (2000).
- Hiraishi et al. [2024] M. Hiraishi, H. Okabe, A. Koda, R. Kadono, T. Muroi, D. Hirai, and Z. Hiroi, Nonmagnetic Ground State in RuO2 Revealed by Muon Spin Rotation, Physical Review Letters 132, 166702 (2024).
- Keßler et al. [2024] P. Keßler, L. Garcia-Gassull, A. Suter, T. Prokscha, Z. Salman, D. Khalyavin, P. Manuel, F. Orlandi, I. I. Mazin, R. Valentı, and S. Moser, Absence of magnetic order in RuO2: insights from SR spectroscopy and neutron diffraction, arXiv:2405.10820 (2024).
- Jovic et al. [2018] V. Jovic, R. J. Koch, S. K. Panda, H. Berger, P. Bugnon, A. Magrez, K. E. Smith, S. Biermann, C. Jozwiak, A. Bostwick, E. Rotenberg, and S. Moser, Dirac nodal lines and flat-band surface state in the functional oxide RuO2, Physical Review B 98, 241101(R) (2018).
- Fedchenko et al. [2024] O. Fedchenko, J. Minár, A. Akashdeep, S. W. D’Souza, D. Vasilyev, O. Tkach, L. Odenbreit, Q. Nguyen, D. Kutnyakhov, N. Wind, L. Wenthaus, M. Scholz, K. Rossnagel, M. Hoesch, M. Aeschlimann, B. Stadtmüller, M. Kläui, G. Schönhense, T. Jungwirth, A. B. Hellenes, G. Jakob, L. Šmejkal, J. Sinova, and H.-J. Elmers, Observation of time-reversal symmetry breaking in the band structure of altermagnetic RuO2, Science Advances 10, eadj4883 (2024).
- Liu et al. [2024] J. Liu, J. Zhan, T. Li, J. Liu, S. Cheng, Y. Shi, L. Deng, M. Zhang, C. Li, J. Ding, Q. Jiang, M. Ye, Z. Liu, Z. Jiang, S. Wang, Q. Li, Y. Xie, Y. Wang, S. Qiao, J. Wen, Y. Sun, and D. Shen, Absence of Altermagnetic Spin Splitting Character in Rutile Oxide RuO2, Physical Review Letters 133, 176401 (2024).
- Lin et al. [2024] Z. Lin, D. Chen, W. Lu, X. Liang, S. Feng, K. Yamagami, J. Osiecki, M. Leandersson, B. Thiagarajan, J. Liu, C. Felser, and J. Ma, Observation of Giant Spin Splitting and -wave Spin Texture in Room Temperature Altermagnet RuO2, arXiv:2402.04995 (2024).
- Huang et al. [1982] Y. Huang, H. Park, and F. H. Pollak, Growth and characterization of RuO2 single crystals, Materials Research Bulletin 17, 1305 (1982).
- Rogers et al. [1969] D. B. Rogers, R. D. Shannon, A. W. Sleight, and J. L. Gillson, Crystal chemistry of metal dioxides with rutile-related structures, Inorganic Chemistry 8, 841 (1969).
- Ryden et al. [1970] W. D. Ryden, A. W. Lawson, and C. C. Sartain, Electrical transport properties of IrO2 and RuO2, Physical Review B 1, 1494 (1970).
- Pawula et al. [2024] F. Pawula, A. Fakih, R. Daou, S. Hébert, N. Mordvinova, O. Lebedev, D. Pelloquin, and A. Maignan, Multiband transport in RuO2, Physical Review B 110, 064432 (2024).
- Kitamura et al. [2022] M. Kitamura, S. Souma, A. Honma, D. Wakabayashi, H. Tanaka, A. Toyoshima, K. Amemiya, T. Kawakami, K. Sugawara, K. Nakayama, K. Yoshimatsu, H. Kumigashira, T. Sato, and K. Horiba, Development of a versatile micro-focused angle-resolved photoemission spectroscopy system with Kirkpatrick–Baez mirror optics, Review of Scientific Instruments 93, 033906 (2022).
- Kresse and Furthmüller [1996] G. Kresse and J. Furthmüller, Efficient iterative schemes for ab initio total-energy calculations using a plane-wave basis set, Physical Review B 54, 11169 (1996).
- Wu et al. [2018] Q. Wu, S. Zhang, H.-F. Song, M. Troyer, and A. A. Soluyanov, WannierTools: An open-source software package for novel topological materials, Computer Physics Communications 224, 405 (2018).
- Mostofi et al. [2008] A. A. Mostofi, J. R. Yates, Y.-S. Lee, I. Souza, D. Vanderbilt, and N. Marzari, wannier90: A tool for obtaining maximally-localised Wannier functions, Computer Physics Communications 178, 685 (2008).
- Smolyanyuk et al. [2024] A. Smolyanyuk, I. I. Mazin, L. Garcia-Gassull, and R. Valentí, Fragility of the magnetic order in the prototypical altermagnet RuO2, Physical Review B 109, 134424 (2024).
- Sun et al. [2017] Y. Sun, Y. Zhang, C.-X. Liu, C. Felser, and B. Yan, Dirac nodal lines and induced spin Hall effect in metallic rutile oxides, Physical Review B 95, 235104 (2017).
- Xu et al. [2019] X. Xu, J. Jiang, W. J. Shi, V. Süß, C. Shekhar, S. C. Sun, Y. J. Chen, S.-K. Mo, C. Felser, B. H. Yan, H. F. Yang, Z. K. Liu, Y. Sun, L. X. Yang, and Y. L. Chen, Strong spin-orbit coupling and Dirac nodal lines in the three-dimensional electronic structure of metallic rutile IrO2, Physical Review B 99, 195106 (2019).
- Nelson et al. [2019] J. N. Nelson, J. P. Ruf, Y. Lee, C. Zeledon, J. K. Kawasaki, S. Moser, C. Jozwiak, E. Rotenberg, A. Bostwick, D. G. Schlom, K. M. Shen, and L. Moreschini, Dirac nodal lines protected against spin-orbit interaction in IrO2, Physical Review Materials 3, 064205 (2019).
- Li et al. [2019a] G. Li, Q. Xu, W. Shi, C. Fu, L. Jiao, M. E. Kamminga, M. Yu, H. Tüysüz, N. Kumar, V. Süß, R. Saha, A. K. Srivastava, S. Wirth, G. Auffermann, J. Gooth, S. Parkin, Y. Sun, E. Liu, and C. Felser, Surface states in bulk single crystal of topological semimetal Co3Sn2S2 toward water oxidation, Science Advances 5, eaaw9867 (2019a).
- Li et al. [2019b] G. Li, C. Fu, W. Shi, L. Jiao, J. Wu, Q. Yang, R. Saha, M. E. Kamminga, A. K. Srivastava, E. Liu, A. N. Yazdani, N. Kumar, J. Zhang, G. R. Blake, X. Liu, M. Fahlman, S. Wirth, G. Auffermann, J. Gooth, S. Parkin, V. Madhavan, X. Feng, Y. Sun, and C. Felser, Dirac Nodal Arc Semimetal PtSn4: An Ideal Platform for Understanding Surface Properties and Catalysis for Hydrogen Evolution, Angewandte Chemie International Edition 58, 13107 (2019b).
Supplemental Material for “Spin-Degenerate Bulk Bands and Topological Surface States of RuO2” by T. Osumi et al.
Supplemental Note 1: Hubbard dependence of the calculated band structure. To clarify the influence of on-site Coulomb interaction energy (Hubbard ) on the calculated Fermi surface (FS) and band structure in the antiferromagnetic (AFM) phase, we compare in Figs. S1(a)–S1(c) and S1(d)–S1(f) the FS in the two-dimensional wave vector and the ARPES-derived band dispersion, respectively, for the (100) surface of RuO2 at at three different values of 1.5, 2.0, and 2.5 eV. We found from our calculations that the AFM phase is energetically more stable than the nonmagnetic (NM) phase at eV, whereas vice versa for eV. Hence, the calculation for eV was carried out for the putative (unstable) AFM phase. One can see in Figs. S1(a)-S1(c) that the calculated FS volume at this slice is systematically reduced upon increasing .
This shrinkage is also seen in the calculated band dispersion in Figs. S1(d)–S1(f). Importantly, none of AFM calculations satisfactorily reproduces the experimental data while the matching between the experiment and calculation becomes gradually better upon reducing . This is reasonable because the reduction of makes the system closer to the NM phase. It is remarked that although the matching between the experiment and calculation becomes the best at eV, the agreement is still poorer than the case of the NM calculation shown in Fig. 2(b) of the main text. Thus, disagreement of the experimental data with the AFM calculation is not due to the incorrect choice of the value, but due to the intrinsic absence of magnetic order in the RuO2 single crystal.
Supplemental Note 2: Surface state at the (100) surface
[…]

### v2

@@ Acknowledgements.
We acknowledge G. Mattoni, H. Matsuki, and C. Sow for their technical supports. We also acknowledge L. Balents and J. N. Hausman for useful comments and discussions. This work was supported by JST-CREST (No. JPMJCR18T1), Grant-in-Aid for Scientific Research (JSPS KAKENHI Grant Numbers JP22H01168, JP21H04435 and JP19H01845), Grant-in-Aid for JSPS Research Fellow (No: JP18J20058), and KEK-PF (Proposal number: 2024S2-001). T. Osumi and A.H. thanks GP-Spin and JSPS for financial support.Appendix A HUBBARD DEPENDENCE OF THE CALCULATED BAND STRUCTURE
To clarify the influence of on-site Coulomb interaction energy (Hubbard ) on the calculated Fermi surface (FS) and band structure in the antiferromagnetic (AFM) phase, we compare in Figs. 6(a)–6(c) and 6(d)–6(f) the FS in the two-dimensional wave vector and the ARPES-derived band dispersion, respectively, for the (100) surface of RuO2 at at three different values of 1.5, 2.0, and 2.5 eV. We found from our calculations that the AFM phase is energetically more stable than the nonmagnetic (NM) phase at eV, whereas vice versa for eV. Hence, the calculation for eV was carried out for the putative (unstable) AFM phase. One can see in Figs. 6(a)–6(c) that the calculated FS volume at this slice is systematically reduced upon increasing .
This shrinkage is also seen in the calculated band dispersion in Figs. 6(d)–6(f). Importantly, none of AFM calculations satisfactorily reproduces the experimental data while the matching between the experiment and calculation becomes gradually better upon reducing . This is reasonable because the reduction of makes the system closer to the NM phase. It is remarked that although the matching between the experiment and calculation becomes the best at eV, the agreement is still poorer than the case of the NM calculation shown in Fig. 2(b) of the main text. Thus, disagreement of the experimental data with the AFM calculation is not due to the incorrect choice of the value, but due to the intrinsic absence of magnetic order in the RuO2 single crystal.
Appendix B SURFACE STATE AT THE (100) SURFACE
To clarify the surface or bulk nature of the observed flat band around the point for the (100) surface presented in Figs. 2(d)–2(i), we carried out -dependent ARPES measurements. Figure 7(a) shows the -dependence of EDC obtained with a normal emission set-up. One can recognize several features dispersing upon variation, which originate from the bulk bands. Corresponding to the flat band, a sharp peak is identified in the vicinity of at = 40 eV. The energy position of this peak is invariant against the variation, supporting its surface nature, while its intensity is strongly modulated by the photoelectron matrix-element effect. These characteristic spectral features are better seen in the ARPES-intensity plot against and in Fig. 7(b). Comparison with the calculated bands in the NM phase signifies a reasonable correspondence of dispersive features between the experiment and calculation, such as an electron band bottomed at 1.5 eV at the point and a holelike band topped at 1.8 eV at the point, confirming its bulk nature. On the other hand, there exist no corresponding bands in the calculation that could account for the peak at , supporting its surface origin.
Appendix C SPIN-RESOLVED ARPES DATA
To reconfirm the absence of altermagnetic band splitting in bulk bands, we have carried out spin-resolved ARPES measurements at CASSIOPÉE beamline in synchrotron SOLEIL. Figures 8(a) and 8(c) show the spin-resolved EDCs at = 40 K measured at a point midway between the and M points for the (100) surface, where the AFM calculation predicts a sizable (more than 0.4 eV) spin splitting. One can immediately recognize no difference between the spin-up and spin-down EDCs along the -axis (we set the and axes parallel to the sample surface at the normal emission with the -axis being along [001]). The observed spin polarization is essentially zero within experimental uncertainty as shown in Figs. 8(b) and 8(d). Here, we estimate the experimental uncertainty, namely the statistical error of the spin polarization. It is written as where is the Sherman function of spin detector and is the number of detection counts [32]. In our measurements, was in the range of 0.2–0.3, and the number of counts for each energy point was 60,000–120,000. These values yield an estimated theoretical of 0.01–0.02, consistent with the experimental standard deviations for the spin polarization shown in Figs. 8(b) and 8(d) [ = 0.006–0.016]. We also found that the polarization along the -axis is also zero (not shown). This supports the absence of altermagnetic band splitting at the (100) surface. We have carried out spin-resolved ARPES measurements also for the (110) surface at the point where the altermagnetic splitting is theoretically predicted in the AFM calculation, but found no clear spin polarization [Figs. 8(e)–8(h)].
Appendix D INFLUENCE OF SOC TO THE CALCULATED BAND STRUCTURE
To clarify the influence of SOC on the calculated band structure of RuO2, we show in Figs. 9(a) and 9(b) a side-by-side comparison of the calculated band dispersions in the NM phase along the XMARX cut () obtained without and with SOC, respectively, overlaid with the corresponding ARPES intensity at the (100) surface. One can see an overall similarity in these calculations, in particular, regarding their (, ) position. A closer look reveals a small band doubling in the calculation with SOC, together with a hybridization gap at the intersection in some of these bands, as highlighted by the band crossing point associated with the DNL along the MA cut in Figs. 9(a) and 9(b). This trend is also seen in the calculation for the XRZ line () shown in Figs. 9(c) and 9(d). The band doubling predicted in the calculation is so small that it is unlikely to be resolved within our experimental accuracy. At the (101) surface, one can see a similar band doubling in the calculations with SOC [Figs. 9(e)–9(h)], whereas the doubling leads to more complex band dispersions at particular cuts like the A cut due to the band hybridization. The calculation with SOC along the A cut shows a relatively good agreement with the ARPES intensity near , implying a finite influence of SOC on the experimental spectral feature.
Appendix E TOPOLOGICAL ANALYSIS OF BULK NODAL LINES
Fermi surfaces in metals possess topological stability when the constituent bands host spin-independent energy degeneracies in the form of nodal points or nodal lines within the BZ [14]. The corresponding topological invariant is given by the Berry phase evaluated either on a closed surface enclosing a nodal point or along a line that pierces a closed nodal loop. When such a degeneracy is accompanied by the band inversion, the invariant acquires a nontrivial, quantized value as a “jump” in the Berry phase, and SSs that converge toward the projected degeneracy points emerge. Many compounds exhibit topological nodal structures and associated SS and are therefore classified as topological semimetals.
We evaluate the topology of RuO2 adopting the same framework as these topological semimetals [10, 17, 13, 12, 25]. RuO2 has a rutile structure (space group ), and several DNLs exist in the bulk BZ, protected by (: space-inversion, : time-reversal) and mirror symmetry. As shown in Fig. 10(a), our bulk band calculations without SOC in the NM phase shown reveal several band crossings in the energy range of approximately 0.3 eV with respect to , originating from band 16 and band 17 (energy bands are sequentially labeled from the lowest one). Along the AM cut, these bands intersect midway between A and M. This Dirac crossing generates an X-shaped nodal line (DNL1) in the 3D BZ at (in units of ), as illustrated by the light-blue curves in Fig. 10(b), which forms the central focus of this study. DNL1 lies on the (110) mirror plane, where bands with opposite mirror eigenvalues cross and keep degeneracy due to the combined mirror and symmetry. In addition to DNL1, the crossings of bands 16 and 17 give rise to several other DNLs (DNL2–7) [Figs. 10(a) and 10(b)], but their characteristics are distinct from those of DNL1. For example, as shown in Fig. 10(a), DNL2/3–7 and DNL4 along high-symmetry cuts are located at approximately -0.3 eV and +0.6 eV, respectively, whereas DNL1 is situated much closer to at eV. While DNL1 forms open nodal lines extending toward neighboring BZs, all the other DNLs exhibit closed-loop structures around the bulk point [Fig. 10(b)].
We evaluate the Zak phase by using equation (1). The Zak phase () corresponds to a invariant, quantized to either 0 or due to the symmetry. In the prototypical nodal-line semimetal Ca3P2 (space group ), the DNL forms a closed loop (i.e. nodal ring), which gives rise to a well-defined region inside the projection of the nodal ring on the surface BZ [13]. Like RuO2, Ca3P2 possesses mirror symmetry with respect to the (001) plane. The nodal ring appears on the = 0 mirror plane forming a single closed nodal loop near the BZ boundary, in which bands keep degeneracy associated with the combined mirror and symmetry. Due to the bulk-edge correspondence, this region hosts nearly dispersion-less drumhead-type SSs, whereas no such states appear outside the ring. When a single DNL is projected onto the surface, the Zak phase takes a value of on one side of the projection and 0 on the other, giving rise to topological SS in the region with . In contrast, when two DNLs are projected onto the surface, the Zak phase becomes 2, which is effectively equivalent to 0.
In the case of RuO2, we found that many DNLs with closed loop structure as in Ca3P2 show which is effectively equivalent to 0 because two DNLs are projected onto the surface. In the right panel of Fig. 10(b), we mapped as a function of the 2D wave vectors in the (100) surface BZ. We found that remains zero (blue region) for the projections of all DNLs except for DNL5, which yields (red region). This is because all DNLs except DNL5 are elongated along the direction and are therefore doubly projected onto the (100) surface for any . In contrast, DNL5 lies in the - plane (i.e. = 0) and is singly projected, resulting in inside its elongated loop. For the (110) surface [bottom panel of Fig. 10(b)], the dog-bone-shaped nodal loop of DNL6 (yellow curves) is singly projected, giving rise to a region with . In the small area where projections of two DNL6 loops overlap (indicated by six white arrows), becomes 0. It should be noted, however, that since all of these DNLs lie relatively far from , the associated topological SS are buried within the bulk bands and cannot be clearly resolved in our ARPES measurements.
We now turn our attention to DNL1 near , which is the primary focus of this study and exhibits topological characteristics distinct from the above case. While the DNL5 forms enclosed loop in the (110) plane (namely ZAM), the DNL1 extends across the BZ, connecting its opposite boundaries without forming a closed loop. Because the DNL1 is open, it does not enclose any finite area, and therefore no well-defined “inside” region can be assigned to the phase. As a result, the Zak phase calculated along the surface-normal [100] direction does not correlate with the presence of the SSs in RuO2. Instead, in the context of mirror symmetry, the mirror eigenvalues of bulk bands are exchanged along (110) plane between the bands 16 and 17. This change in the mirror representation acts as an effective topological boundary within the BZ from which the SS emerges. Thus, the existence of the SS is not governed by the one-dimensional Zak phase like in Ca3P2, but by the band connectivity protected by the mirror and symmetries associated with the open nodal lines [61]. In this sense, the SSs on the (100) surface of RuO2 relevant to DNL1 reflect the open nodal-line topology which is not captured by a phase, in contrast to conventional drumhead SSs defined by a region.
Analogous to the case of a nodal line protected by mirror and symmetries, each crossing for DNL1 causes a jump in the Zak phase due to the inversion of the occupied-band mirror eigenvalues. This is confirmed by our calculation of Wilson loop encircling a single DNL1 [see the three red circles in Fig. 5(a) of the main text] that consistently yields a value of , demonstrating the intrinsic topological nature of DNL1. However, since the two mirror planes, (110) and (10), are symmetry-related and induce opposite phase inversions, the two shifts cancel each other, resulting in a net Zak phase of = 0 along the [100] and [110] directions. Thus, no well-defined region can be defined, even though SSs can still emerge around the projections of the nodal lines protected by mirror and symmetries. Namely, the dispersion of SSs of RuO2 on the (100) surface exhibits a drumhead-like shape but lacks a corresponding topological character. Consequently, on both the (100) and (110) surfaces where DNL1 is doubly projected, two topological SS (assuming spin-degenerate bands) are expected to emerge from the Dirac node when SOC is neglected, as illustrated in the top panel of Fig. 5(c). Although SOC open a finite gap along the former DNL, surface-related states remain in the corresponding -region, reflecting the residual band connectivity that exists without SOC, as shown in the bottom panel of Fig. 5(c).
Appendix F SLAB CALCULATIONS
To clarify the characteristics of the observed SSs in more detail, we have carried out slab calculations in the NM phase for the (100), (110), and (101) surfaces by assuming the oxygen-deficient termination as shown in Figs. 11(a)–(c) [corresponding BZs are also shown in Figs. 11(d)–(f)]. We found that the oxygen-deficient termination reproduces the experimental results relatively well, as detailed below.
F.1 (100) surface
Figure 12(a) shows the calculated band dispersion along high-symmetry lines in the surface BZ. In this calculation, SOC was intentionally neglected in order to highlight the connection between the DNL and the topological SS. Two SSs, labeled SS1 and SS2, are connected to DNL1 along the cut, as expected from the schematic in Fig. 5(c). Consistent with the discussion based on the mirror eigenvalues in previous section (Appendix E), the calculated flat SSs on the (100) surface bridges the projections of nodal lines, instead of filling an enclosed area. Although the DNL1 cannot be characterized by a phase, the resulting surface bands (SS1 and SS2) exhibit a drumhead-like flat dispersion in certain regions of the BZ (e.g. along ) and can be interpreted as a symmetry-protected SSs associated with the nodal-line topology.
Next we have included SOC in our slab calculations. As a result, we obtained a stable solution with a magnetic moment (-0.4, -0.4, -0.4) in the component induced in the surface oxygen atoms. In Figs. 12(b) and 12(c), we present two cases; one in which the total surface magnetic moment was artificially constrained to zero, and another in which no such constraint was imposed. In Fig. 12(b), one can directly identify the influence of SOC. The overall energy dispersion of SS1 and SS2 remain nearly unchanged, except for a very small spin splitting. This indicates that the effect of SOC on the SS dispersion is relatively weak.
In both Fig. 12(a) and 12(b), the flat band appears along the line, as in the experiment, but is located at 0.45 eV above , in contrast to our ARPES data which show the flat band pinned almost at . This discrepancy is resolved in the most stable solution, shown in Fig. 12(c) in which a band-dependent exchange splitting of the SS can be seen; notably, the flat band around the point now appears near , corresponding to the lower branch of the spin-split SS1 (SS1). Although the validity of possible surface ferromagnetism remains to be clarified, this slab calculation exhibits much better agreement with the ARPES data, as detailed further below.
F.2 (110)/(101) Surfaces
To further validate the topological character of DNL1, we have investigated bulk-edge correspondence in other surface orientations, (110) and (101). In the (110) case, DNL1 extends along a momentum direction perpendicular to the surface [Fig. 10(b)]. This causes DNL1 to be buried within the projection of bulk bands, as illustrated in Fig. 13(b). Specifically, a band crossing point at +0.2 eV is observed midway between and , originating from the crossing point of DNL1 along the line in bulk BZ [Fig. 5(d) in main text]. Similarly, another band crossing point appears near , close to the point, originating from the crossing point of DNL1 along the MA line in the bulk BZ. Although the correspondence between the SSs and DNL1 is less clear in the (110) orientation compared to the (100) case, two surface bands can still be identified, extending from the DNL1 crossing point (along line) toward . These bands exhibit downward dispersions and form flat features at and -0.2 eV. Importantly, the flat band at shows good correspondence with that observed in the ARPES data on the (110) surface [yellow circles in Fig. 13(c)].
[…]